OpenStudy (kevin):

1/2 + 1/6 + 1/12 + 1/20 +… + 1/9900 =

OpenStudy (kevin):

@mathmate @brainzonly

OpenStudy (mathmate):

\(\Large \sum_{i=1}^{99} \frac{1}{i(i+1)}\), right?

OpenStudy (agent0smith):

Just need to find a sequence that describes 2, 6, 12, 20, 30, 42 +4, +6, +8, +10, +12 +2, +2, +2 By eyeballing it, it's \[\large n^2 + n\] and for 9900 = n^2+n, the last value is n=99 So your series is \[\large \sum_{n=1}^{n=99} \frac{ 1 }{ n^2 + n }\]

OpenStudy (mathmate):

If so, then use \(\Large \frac{1}{i(i+1)}\equiv\frac{1}{i}-\frac{1}{i+1}\) and sum the truncated harmonic series.

OpenStudy (kevin):

that's very complicated :/ Can you explain me further or use easier way?

OpenStudy (mathmate):

hint: read about telescopic series.

OpenStudy (kevin):

ok

OpenStudy (agent0smith):

Which part don't you understand?

OpenStudy (brainzonly):

You look at the denominator, and you will find that 2,6,12 ... ... Can be obtained by multiplying two consecutive natural numbers. You can type turns into the original 1/1*2+1/2*3+1/3*4+1/4*5+......+1/99*100=1-1/2+1/2-1/3+1/3-1/4+1/4-1/5+......+1/99-1/100 Parts are found in the middle of all the simplicity of the AH? Above I have no parentheses, bored, you can read it. And =1-1/100=99/100 The answer is 99/100. Later this topic to learn how to analyze and question you must be characteristic of the flexibility of thinking to analyze and understand, of course, long after will naturally think of the method.

OpenStudy (kevin):

@agent0smith how do you know 9900 = n^2+n, the last value is n=99

OpenStudy (agent0smith):

By solving for n.\[\large 9900 = n^2+n\]\[\large 0 = n^2+n - 9900 = (n+100)(n-99)\]

OpenStudy (kevin):

Owhh.. I see

OpenStudy (kevin):

Then?

OpenStudy (agent0smith):

See @brainzonly's method, it's an easy way to find the sum.

OpenStudy (kevin):

@brainzonly How do you get it's same with 1 - 1/100 ?

OpenStudy (mathmate):

The sum breaks down to \(\Large \sum_{n=1}^{100} \frac{1}{n}-\frac{1}{n+1}\) =(1/1-1/2)+(1/2-1/3)+(1/3-1/4)....(1/98-1/99)+(1/99-1/100) [note that the middle terms cancel out, in a telescopic series] =1-1/100 =99/100

OpenStudy (kevin):

I got it. Thx guys!

OpenStudy (agent0smith):

Using what @mathmale and plugging in the first few values of i=1, 2, 3 should show it to you... \[\large \frac{1}{i}-\frac{1}{i+1} =\left( \frac{1}{1}-\frac{1}{1+1} \right) + \left( \frac{1}{2}-\frac{1}{2+1} \right) +\left( \frac{1}{3}-\frac{1}{3+1} \right) +...\] \[\large \frac{1}{i}-\frac{1}{i+1} =\left( \frac{1}{1}-\frac{1}{2} \right) + \left( \frac{1}{2}-\frac{1}{3} \right) +\left( \frac{1}{3}-\frac{1}{4} \right) +...\]

OpenStudy (agent0smith):

@mathmate not mathmale

OpenStudy (mathmate):

@agent0smith :) we're so confusing!

OpenStudy (brainzonly):

That's why everybody likes ME. :D @mathmate is probably smarter though.